# and the antiproton puzzle

### v1

@@ Abstract
We analyze joint factorial cumulants of protons and antiprotons in relativistic heavy-ion collisions and point out that they obey the scaling $\hat{C}_{nm}^{p,\bar{p}}\propto\left\langle N_{p}\right\rangle^{n}\left\langle N_{\bar{p}}\right\rangle^{m}$ as a function of acceptance when only long-range correlations are present in the system, such as global baryon conservation and volume fluctuations. This hypothesis can be directly tested experimentally without the need for corrections for volume fluctuations. We show that if correlations among protons and antiprotons are driven by global baryon conservation and volume fluctuations only, the equality $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}=\hat{C}_{2}^{\bar{p}}/\left\langle N_{\bar{p}}\right\rangle^{2}$ holds for large systems created in central collisions. We point out that the experimental data of the STAR Collaboration from phase I of RHIC beam energy scan are approximately consistent with the scaling $\hat{C}_{nm}^{p,\bar{p}}\propto\left\langle N_{p}\right\rangle^{n}\left\langle N_{\bar{p}}\right\rangle^{m}$, but the normalized antiproton correlations are stronger than that of protons, $-\hat{C}_{2}^{\bar{p}}/\left\langle N_{\bar{p}}\right\rangle^{2}>-\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$, indicating that global baryon conservation and volume fluctuations alone cannot explain the data. We also discuss high-order factorial cumulants which can be measured with sufficient precision within phase II of RHIC-BES.
@@ I Introduction
Exploring the structure of the QCD phase diagram and searching for a possible phase transition and an associated critical point is one of the major goals in strong interaction research (see Bzdak et al. (2020) for a review). One of the key observables are fluctuations of globally conserved charges in a subsystem of the fireball created in high-energy nucleus-nucleus collisions, most prominently (net) baryon number fluctuations. Fluctuation observables have the advantage that they are accessible both to experiment and theory, such as Lattice QCD. However, care needs to be taken to make comparisons between theory and experiment. Theoretical calculations typically deal with a system with fixed volume in the grand-canonical ensemble, where the charges are conserved only on the average. In addition, they calculate fluctuations of the baryon number. In contrast, for the system studied in heavy-ion collision experiment, the charges are globally conserved and, even for the best centrality selection the volume of the system fluctuates. Also, experiments can only measure the fluctuation of protons, and possibly lambdas and light nuclei, but not those of all baryons. Therefore, even for a system without any phase transition, the fluctuation observables extracted in the experiment differ from the naive theoretical expectation, and appropriate corrections need to be applied for a meaningful comparison.
@@
Here, we discuss such observables, namely the so-called reduced correlation coefficients, which are the factorial cumulants scaled with proper powers of the mean particle number and have earlier been introduced and discussed in Bzdak et al. (2017a, b); Bzdak and Koch (2017); Barej and Bzdak (2020). We will show that these reduced correlation coefficients should be independent of the acceptance if the observed fluctuations arise solely from a combination of statistical fluctuations, volume fluctuations, and global baryon number conservation. We will further show that the second-order reduced correlation coefficient should be (to a very good approximation) the same for protons and antiprotons. We will then analyze the data of the STAR collaboration from the first phase of the RHIC beam energy scan (BES-I) Abdallah et al. (2021) as to which extent the fluctuation data are consistent with this simple baseline or if there are indications for additional, potentially critical dynamics.
@@ II.3 Global baryon conservation
$P(N_{B},N_{\bar{B}})\propto\frac{\left\langle N_{B}\right\rangle}{N_{B}!}\frac{\left\langle N_{\bar{B}}\right\rangle}{N_{\bar{B}}!}\delta_{B,N_{B}-N_{\bar{B}}},$ (11)
from which one can derive the following cumulant generating function
@@
$=\ln\left[e^{B\frac{t_{B}-t_{\bar{B}}}{2}}\frac{I_{B}(2ze^{\frac{t_{B}+t_{\bar{B}}}{2}})}{I_{B}(2z)}\right].$ (12)
@@ II.4 Volume fluctuations
[…]
Here $P_{V}(N_{B},N_{\bar{B}})$ is the baryon and anti-baryon distribution at fixed volume and $\left\langle\ldots\right\rangle_{V}$ indicates averaging at a fixed volume. Thus we get for the cumulant generating function $G_{\tilde{N}_{B},\tilde{N}_{\bar{B}}}$ for cumulants $\tilde{\kappa}_{n,m}$ affected by volume fluctuations
@@ III.1 BES-I data
The STAR Collaboration has presented measurements of ordinary and factorial cumulants for protons and antiprotons in Au-Au collisions at $\sqrt{s_{\rm NN}}=7.7-200$ GeV in Adam et al. (2021); Abdallah et al. (2021) within phase I of RHIC beam energy scan. The absolute values of the cumulants as well as the scaled factorial cumulants such as $\hat{C}_{n}^{p}/\left\langle N_{p}\right\rangle$ or $\hat{C}_{n}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle$ were published in Abdallah et al. (2021), but not the corresponding couplings $\hat{c}_{n}^{p}=\hat{C}_{n}^{p}/\left\langle N_{p}\right\rangle^{n}$ etc. Figure 1 shows the rapidity cut dependence of the couplings for protons, $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$ (blue), and antiprotons, $\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}$ (black). The experimental data indicate that the scalings $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}\propto\rm const$ and $\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}\propto\rm const$ approximately hold, with some indications for additional negative two-proton correlations at small $y_{\rm cut}$ at $\sqrt{s_{\rm NN}}=14.5-27$ GeV. Interestingly, one can clearly see that anticorrelations of antiprotons are notably stronger than those of the protons ones at energies below $\sqrt{s_{\rm NN}}\leq 62.4$ GeV, where $-\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}>-\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$.
@@ III.2 Single fireball model
Thus, the description of antiproton cumulants within a hydrodynamics-based single fireball picture remains an open challenge that needs further investigation. One possible avenue is the incorporation of re-scattering in the hadronic phase and baryon annihilation, the latter are expected to affect antiprotons significantly Becattini et al. (2017) and leave an imprint on their cumulants Savchuk et al. (2022).
@@ III.3 Two-source model
The protons contain contributions from both stopped and produced matter, $\left\langle N_{p}\right\rangle=\left\langle N_{p}\right\rangle_{\rm stopped}+\left\langle N_{p}\right\rangle_{\rm prod}$. Since the produced matter is net-baryon free, one has $\left\langle N_{p}\right\rangle_{\rm prod}=\left\langle N_{\bar{p}}\right\rangle$ and thus $\left\langle N_{p}\right\rangle_{\rm stopped}=\left\langle N_{p}\right\rangle-\left\langle N_{\bar{p}}\right\rangle$. Here we model the stopped protons by a binomial distribution with Bernoulli probability of $p=\left\langle p\right\rangle_{\rm stopped}/N_{part}$. Since the two sources are assumed to be independent, the cumulants simply add. Thus, assuming a fixed number of participants, the second-order proton factorial cumulant reads
@@
The two-source model thus qualitatively describes the splitting between protons and antiprotons. However, as one can see from Fig. 1, this splitting is overestimated at energies $\sqrt{s_{\rm NN}}\leq 39$ GeV. This may indicate that the two sources are not completely independent but also not fully thermalized, as the latter scenario would correspond to the one-component model where there is no splitting. It should be noted that produced matter is assumed to be thermalized here, such that the cumulants of produced protons and antiprotons are described by the ideal gas in the canonical ensemble. In the opposite case Tomasik et al. (2019), one may assume that produced $B\bar{B}$ pairs are uncorrelated, and their distribution follows the Poisson statistics. In this case, factorial cumulants of antiprotons are zero starting from the 2nd order, in contrast to the data. The experimental data, which lie in-between these two scenarios at $\sqrt{s_{\rm NN}}\lesssim 39$ GeV may thus suggest also incomplete equilibration of produced matter.
We note in the case of a two-source model, volume fluctuations become more tricky, as here, one may in principle have to consider two volumes, one for the participant system and one for the produced system. For instance, if participant fluctuations are negligible compared to fluctuations of the volume for the produced matter, i.e. of $N_{\rm pairs}^{B\bar{B}}$, one will expect volume fluctuations to affect antiproton $\hat{c}_{2}^{\bar{p}}$ more strongly, and such mechanism could help in describing the experimental data.
@@ III.4 Short-range repulsion
These results indicate that non-critical correlations such as excluded volume would be very challenging to see through the analysis of $c_{2}^{p}=\rm const$ scaling and its possible violation, at least with second order cumulants. On the other hand, critical fluctuations are expected to yield a more pronounced acceptance dependence of factorial cumulants Ling and Stephanov (2016), especially beyond the second order.
@@ III.5 LHC energies
The ALICE Collaboration has performed measurements of (net) proton number cumulants in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=2.76$ TeV Acharya et al. (2020) and $\sqrt{s_{\rm NN}}=5.02$ TeV Acharya et al. (2023). The main focus has been on the variance of net proton number normalized by the Skellam (Poisson statistics) baseline, $R_{1}=\frac{\kappa_{2}(p-\bar{p})}{\left\langle p+\bar{p}\right\rangle}$ Acharya et al. (2020). In terms of factorial cumulants, this quantity reads
@@
The ALICE Collaboration has presented pseudorapidity acceptance dependence of $R_{1}$ in Acharya et al. (2020, 2023), but unfortunately not that of $\left\langle N_{p}+N_{\bar{p}}\right\rangle$. This makes it impossible to verify the scaling of $r_{1}$ with acceptance using available published data. In Acharya et al. (2020) the values of $\left\langle N_{p}+N_{\bar{p}}\right\rangle$ for the largest acceptance $|\eta|<0.8$ have been published, for $\sqrt{s_{\rm NN}}=2.76$ TeV Pb-Pb collisions. With $R_{1}=0.971\pm 0.015$ and $\left\langle p+\bar{p}\right\rangle=36.49\pm 0.65$, one can estimate $r_{1}=(-0.77\pm 0.40)\cdot 10^{-3}$ and thus
@@ IV Discussion and summary
  • However, the data show a clear difference between the reduced correlators of protons and anti-protons, except for the highest energy of $\sqrt{s_{\rm NN}}=200$ GeV. In particular the reduced correlator of anti-protons is not described by the non-critical hydrodynamics baseline or in other words we do have an anti-proton puzzle. Here we considered the extreme scenario of a two source model with separate sources for produced and stopped protons. While this model gives the observed trend it clearly over-predicts the effect. We have not considered proton – anti-proton annihilation which at the lowest energies would lead to different results for protons and anti-protons. However, we note that the difference is still present at a beam energy of $\sqrt{s}=62.4\,\rm GeV$ where annihilation effects should be very similar for protons and anti-protons.
  • We have shown that the effect of volume fluctuations also leads to flat acceptance dependence of the reduced correlators. Therefore, it would be interesting if the STAR data without the so-called Centrality Bin Width Corrections (CBWC) Luo et al. (2013) also exhibited a flat behavior. If not, this implies that the CBWC procedure does more than just suppress volume fluctuations.
@@
  • While our two source model quantitatively over-predicts the difference between the reduced correlators between protons and anti-protons, one could argue that the STAR data suggest that we are dealing with more than one single source. However, before such a conclusion can be reached, the effect of baryon annihilation needs to be studied in detail.

### v2

@@ Abstract
We analyze joint factorial cumulants of protons and antiprotons in relativistic heavy-ion collisions and point out that they obey the scaling $\hat{C}_{nm}^{p,\bar{p}}\propto\left\langle N_{p}\right\rangle^{n}\left\langle N_{\bar{p}}\right\rangle^{m}$ as a function of acceptance when only long-range correlations are present in the system, such as global baryon conservation and volume fluctuations. This hypothesis can be directly tested experimentally without the need for corrections for volume fluctuations. We show that if correlations among protons and antiprotons are driven by global baryon conservation and volume fluctuations only, the equality $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}=\hat{C}_{2}^{\bar{p}}/\left\langle N_{\bar{p}}\right\rangle^{2}$ holds for large systems created in central collisions. We point out that the experimental data of the STAR Collaboration from phase I of RHIC beam energy scan are approximately consistent with the scaling $\hat{C}_{nm}^{p,\bar{p}}\propto\left\langle N_{p}\right\rangle^{n}\left\langle N_{\bar{p}}\right\rangle^{m}$, but the normalized antiproton correlations are stronger than that of protons, $\hat{C}_{2}^{\bar{p}}/\left\langle N_{\bar{p}}\right\rangle^{2}>-\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$. Existing theoretical baselines, based on global baryon conservation and volume fluctuations, cannot explain the data, to which we refer as the antiproton puzzle. We also discuss high-order factorial cumulants which can be measured with sufficient precision within phase II of RHIC-BES.
@@ I Introduction
Exploring the structure of the QCD phase diagram and searching for a possible phase transition and an associated critical point is one of the major goals in strong interaction research (see Bzdak et al. (2020) for a review). One of the key observables are fluctuations of globally conserved charges in a subsystem of the fireball created in high-energy nucleus-nucleus collisions, most prominently (net) baryon number fluctuations. Fluctuation observables have the advantage that they are accessible both to experiment and theory, such as Lattice QCD. However, care needs to be taken to make comparisons between theory and experiment. Theoretical calculations typically deal with a system with fixed volume in the grand-canonical ensemble, where the charges are conserved only on the average. In addition, they calculate fluctuations of the baryon number. In contrast, for the system studied in heavy-ion collision experiments, the charges are globally conserved, and even for the best centrality selection, the volume of the system fluctuates. Also, experiments can only measure the fluctuation of protons, and possibly lambdas and light nuclei, but not those of all baryons. Therefore, even for a system without any phase transition, the fluctuation observables extracted in the experiment differ from the naive theoretical expectation, and appropriate corrections need to be applied for a meaningful comparison.
@@
Here, we discuss such observables, namely the so-called reduced correlation coefficients $\hat{c}_{n}=\hat{C}_{n}/\left\langle N\right\rangle^{n}$, which are the factorial cumulants scaled with proper powers of the mean particle number and have earlier been introduced and discussed in Bzdak et al. (2017a, b); Bzdak and Koch (2017); Barej and Bzdak (2020). We will show that these reduced correlation coefficients should be independent of the acceptance if the observed fluctuations arise solely from a combination of statistical fluctuations, volume fluctuations, and global baryon number conservation. We will further show that the second-order reduced correlation coefficient should be (to a very good approximation) the same for protons and antiprotons.
We then analyze the data of the STAR collaboration from the first phase of the RHIC beam energy scan (BES-I) Abdallah et al. (2021) to see if the fluctuation data are consistent with this simple baseline or if there are indications for additional, potentially critical dynamics. Given the statistics limitations of the BES-I data, the analysis here is restricted to second-order coefficients. Although the BES-I data do not reveal a significant variation in the acceptance dependence, one does observe a remarkably strong splitting in the strength of the second-order reduced correlation coefficient between protons and antiprotons. While the protons can be reasonably well described by global baryon conservation-based baselines, the strength of antiproton anticorrelation is significantly underestimated by the available baselines, to which we refer to as the antiproton puzzle. We then discuss the possible mechanisms that may be responsible for the observed splitting, such as the possible incomplete equilibration between stopped and produced matter.
@@ II.3 Global baryon conservation
$P(N_{B},N_{\bar{B}})\propto\frac{\left\langle N_{B}\right\rangle^{N_{B}}}{N_{B}!}\frac{\left\langle N_{\bar{B}}\right\rangle^{N_{\bar{B}}}}{N_{\bar{B}}!}\delta_{B,N_{B}-N_{\bar{B}}},$ (11)
from which one can derive the following cumulant generating function Bzdak et al. (2013)
@@
$=\ln\left[\exp\left(B\,\frac{t_{B}-t_{\bar{B}}}{2}\right)\frac{I_{B}(2ze^{\frac{t_{B}+t_{\bar{B}}}{2}})}{I_{B}(2z)}\right].$ (12)
@@ II.4 Volume fluctuations
$M_{\tilde{N}_{B},\tilde{N}_{\bar{B}}}=\int dV\rho_{V}(V)\left\langle e^{t_{B}N_{B}+t_{\bar{B}}N_{\bar{B}}}\right\rangle_{V}$ (26)
Here $P_{V}(N_{B},N_{\bar{B}})$ is the baryon and anti-baryon distribution at fixed volume and $\left\langle\ldots\right\rangle_{V}$ indicates averaging at a fixed volume. We thus obtain the cumulant generating function $G_{\tilde{N}_{B},\tilde{N}_{\bar{B}}}$ for cumulants $\tilde{\kappa}_{n,m}$ affected by volume fluctuations
@@ III.1 BES-I data
The STAR Collaboration has presented measurements of ordinary and factorial cumulants for protons and antiprotons in Au-Au collisions at $\sqrt{s_{\rm NN}}=7.7-200$ GeV in Adam et al. (2021); Abdallah et al. (2021) within phase I of RHIC beam energy scan. The absolute values of the cumulants as well as the scaled factorial cumulants such as $\hat{C}_{n}^{p}/\left\langle N_{p}\right\rangle$ or $\hat{C}_{n}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle$ were published in Abdallah et al. (2021), but not the corresponding couplings $\hat{c}_{n}^{p}=\hat{C}_{n}^{p}/\left\langle N_{p}\right\rangle^{n}$ etc. Figure 1 shows the rapidity cut dependence of the couplings for protons, $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$ (blue), and antiprotons, $\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}$ (black). The experimental data indicate that the scalings $\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}\propto\rm const$ and $\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}\propto\rm const$ approximately hold, with some indications for additional negative two-proton correlations at small $y_{\rm cut}$ at $\sqrt{s_{\rm NN}}=14.5-27$ GeV. Interestingly, one can clearly see that anticorrelations of antiprotons are notably stronger than those of the protons at energies below $\sqrt{s_{\rm NN}}\leq 62.4$ GeV, where $\hat{C}_{2}^{{\bar{p}}}/\left\langle N_{\bar{p}}\right\rangle^{2}>-\hat{C}_{2}^{p}/\left\langle N_{p}\right\rangle^{2}$.
@@ III.2 Single fireball model
Thus, the description of antiproton cumulants within a hydrodynamics-based single fireball picture remains an open challenge – the antiproton puzzle – that needs further investigation. One possible avenue is the incorporation of re-scattering in the hadronic phase and baryon annihilation, the latter are expected to affect antiprotons significantly Becattini et al. (2017) and leave an imprint on their cumulants Savchuk et al. (2022).
@@ III.3 Two-source model
The protons contain contributions from both stopped and produced matter, $\left\langle N_{p}\right\rangle=\left\langle N_{p}\right\rangle_{\rm stopped}+\left\langle N_{p}\right\rangle_{\rm prod}$. Since the produced matter is net-baryon free, one has $\left\langle N_{p}\right\rangle_{\rm prod}=\left\langle N_{\bar{p}}\right\rangle$ and thus $\left\langle N_{p}\right\rangle_{\rm stopped}=\left\langle N_{p}\right\rangle-\left\langle N_{\bar{p}}\right\rangle$. Here we model the stopped protons by a binomial distribution with Bernoulli probability of $p=\left\langle N_{p}\right\rangle_{\rm stopped}/N_{\rm part}$. Since the two sources are assumed to be independent, the cumulants simply add. Thus, assuming a fixed number of participants, the second-order proton factorial cumulant reads
@@
The two-source model thus qualitatively describes the splitting between protons and antiprotons. However, as one can see from Fig. 1, this splitting is overestimated at energies $\sqrt{s_{\rm NN}}\leq 39$ GeV. This may indicate that the two sources are not completely independent but also not fully thermalized, as the latter scenario would correspond to the one-component model where there is no splitting. It should be noted that produced matter is assumed to be thermalized here, such that the cumulants of produced protons and antiprotons are described by the ideal gas in the canonical ensemble. In the opposite case Tomasik et al. (2019), one may assume that produced $B\bar{B}$ pairs are uncorrelated, and their distribution follows the Poisson statistics. In this case, factorial cumulants of antiprotons are zero starting from the second order, in contrast to the data. The experimental data, which lie in-between these two scenarios at $\sqrt{s_{\rm NN}}\lesssim 39$ GeV, may thus suggest an incomplete equilibration of produced matter.
We note in the case of a two-source model, volume fluctuations become more tricky, as here, one may, in principle, have to consider two volumes, one for the participant system and one for the produced system. For instance, if participant fluctuations are negligible compared to fluctuations of the volume for the produced matter, i.e. of $N_{\rm pairs}^{B\bar{B}}$, one will expect volume fluctuations to affect antiproton $\hat{c}_{2}^{\bar{p}}$ more strongly, and such mechanism could help in describing the experimental data.
@@ III.4 Short-range repulsion
These results indicate that non-critical correlations such as excluded volume would be very challenging to identify through the analysis of $c_{2}^{p}=\rm const$ scaling and its possible violation, at least with second order cumulants. On the other hand, critical fluctuations are expected to yield a more pronounced acceptance dependence of factorial cumulants Ling and Stephanov (2016), especially beyond the second order.
@@ III.5 LHC energies
The ALICE Collaboration has performed measurements of (net) proton number cumulants in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=2.76$ TeV Acharya et al. (2020) and $\sqrt{s_{\rm NN}}=5.02$ TeV Acharya et al. (2023). The main focus has been on the variance of net proton number normalized by the Skellam (Poisson statistics) baseline, $R_{1}=\frac{\kappa_{2}(N_{p}-N_{\bar{p}})}{\left\langle N_{p}+N_{\bar{p}}\right\rangle}$ Acharya et al. (2020). In terms of factorial cumulants, this quantity reads
@@
The ALICE Collaboration has presented pseudorapidity acceptance dependence of $R_{1}$ in Acharya et al. (2020, 2023), but unfortunately not that of $\left\langle N_{p}+N_{\bar{p}}\right\rangle$. This makes it impossible to verify the scaling of $r_{1}$ with acceptance using available published data. In Acharya et al. (2020) the values of $\left\langle N_{p}+N_{\bar{p}}\right\rangle$ for the largest acceptance $|\eta|<0.8$ have been published, for $\sqrt{s_{\rm NN}}=2.76$ TeV Pb-Pb collisions. With $R_{1}=0.971\pm 0.015$ and $\left\langle N_{p}+N_{\bar{p}}\right\rangle=36.49\pm 0.65$, one can estimate $r_{1}=(-0.77\pm 0.40)\cdot 10^{-3}$ and thus
@@ IV Discussion and summary
  • However, the data show a clear difference between the reduced correlators of protons and anti-protons, except for the highest energy of $\sqrt{s_{\rm NN}}=200$ GeV. In particular the reduced correlator of anti-protons is not described by the non-critical hydrodynamics baseline or in other words we do have an antiproton puzzle. Here we considered the extreme scenario of a two source model with separate sources for produced and stopped protons. While this model gives the observed trend it clearly over-predicts the effect. We have not considered proton – anti-proton annihilation which at the lowest energies would lead to different results for protons and anti-protons. However, we note that the difference is still present at a beam energy of $\sqrt{s}=62.4\,\rm GeV$ where annihilation effects should be similar for protons and anti-protons.
  • We have shown that the effect of volume fluctuations also leads to flat acceptance dependence of the reduced correlators. Therefore, it would be interesting if the STAR data without the so-called Centrality Bin Width Correction (CBWC) Luo et al. (2013) also exhibited a flat behavior. If not, this implies that the CBWC procedure does more than just suppress volume fluctuations.
@@
  • While our two-source model quantitatively over-predicts the difference between the reduced correlators between protons and anti-protons, one could argue that the STAR data suggest that we are dealing with more than one single source. However, before such a conclusion can be reached, the effect of baryon annihilation needs to be studied in detail. Additionally, the effect of local baryon conservation at non-zero baryon density needs to be considered. Measurements of balance functions, which can be connected to the two-particle correlation coefficients studied here may be particularly useful for clarifying the situation Bass et al. (2000); Pruneau et al. (2023).